OpenStudy (pounce0129):

Suppose y varies directly as x. Write a direct variation equation that relates x and y. Then solve. if y=3/4 when x=24, find y when x= 12

jhonyy9 (jhonyy9):

so us the rule of 3 simple and will get when x=24 ----- y=3/4 x=12 ---- y=? so you see that 12 is 24 divided by 2 so than you get 24/2=12 ----- 3/4/2 = 3/(4*2) =3/8

OpenStudy (pounce0129):

i used cross multiplying and i got 0.75

jhonyy9 (jhonyy9):

so 0,75 for what ?

jhonyy9 (jhonyy9):

ok for y

OpenStudy (pounce0129):

y=0.75

OpenStudy (pounce0129):

hello?

jhonyy9 (jhonyy9):

so if you write 24x=(3/4)y 12x=ßy ---------- cross multiplie and calcule ß than will get 24x*ßy =(3/4)y*12x divide both sides by 12x and will get 2ßy=(3/4)y multiplie both sides by 4 8ßy=4(3/4)y on the right side simplifie by 4 8ßy=3y divide both sides by y 8ß=3 divide both sides by 8 ß=3/8 so than resulted that for y=3/8 will be x=12 ok ? do you understand it sure right ?

OpenStudy (pounce0129):

what does ß mean? and yes i get it thanks :)

jhonyy9 (jhonyy9):

so ß mean here the number what need multiplie y for you get x=12

jhonyy9 (jhonyy9):

ok ?

OpenStudy (pounce0129):

thanks

jhonyy9 (jhonyy9):

yw was my pleasure good luck bye